Sara purchases a 3-year bond with a par value of $100, 4% annual coupons, and redeemable at $104. The annual effective yield rate is 6% a) Find the bond purchase price P- b) The bond is sold at a A. premium B. discount c) Fill in the the bond amortization schedule Amount for Period Coupon Interest Earned Accumulation o Book Value Discount 0 Total Note that ΣΡ, C-P-The Amount of Premium. The principal repaid is called the amount for accumulation of discount. It is also called the write up because it increases the book value of the bond from P to
Answer #1

Annual Coupon Bond = 4% x $100 = $ 4 ( Paymable for 3 years at the end of each year)

Interest Rate i = 6 %

Also 104 $ is received at the end of year 3.

a) Purchase Price P = 4 x (1 - 1.06-3)/.06 + 104 x 1.06-3 = 10.69 + 87.32 = 98.01

----------------------------------------------------------------------------------------------------------------------------------------------------

b) Since the purchase price 98.01 < face value of 100, hence bond is sold at discount.

----------------------------------------------------------------------------------------------------------------------------------------------------

c)

Period Coupon Interest Earned Amount for accumulation of Discount Book Value
t Fr i(t) = iB(t-1) P(t) = I(t) - Fr B(t) = B(t-1) + P(t)
0                             -                               -                               -                            98.01
1                        4.00                        5.88                        1.88                          99.89
2                        4.00                        5.99                        1.99                       101.88
3                        4.00                        6.11                        2.11                       104.00

---------------------------------------------------------------------------------------------------------------------------------------------------------